If cos A =5/13 than find the value of tan AHhnnnnfgubvgvnm

Grade:10

1 Answers

Susmita
5 years ago
cosA=5/13
sinA=root(1-cos2A)=root(1-25/169)=root(144/169)=12/13
so tanA=sinA/cisA=12/5
Please approve if it helps.thank you.